### v1

# Cutwidth Bounds via Vertex Partitions

Abstract
We study the cutwidth measure on graphs and ways to bound the cutwidth of a graph by partitioning its vertices. We consider bounds expressed as a function of two quantities: on the one hand, the maximal cutwidth $x$ of the subgraphs induced by the classes of the partition, and on the other hand, the cutwidth $y$ of the quotient multigraph obtained by merging each class to a single vertex. We consider in particular decomposition of directed graphs into strongly connected components (SCCs): in this case, $x$ is the maximal cutwidth of an SCC, and $y$ is the cutwidth of the directed acyclic condensation multigraph.
We show that the cutwidth of a graph is always in $O(x+y)$, specifically it can be upper bounded by $1.5x+y$. We also show a lower bound justifying that the constant $1.5$ cannot be improved in general
1 Introduction
The measure of cutwidth (see cutwidth) is a parameter on undirected graphs which intuitively measures the best performance of a linear ordering of the vertices in terms of how many edges are “cut”. Cutwidth is related to pathwidth: it is always as least as high as the pathwidth, and the two parameters coincide up to constant factors when the maximal degree of the graph is bounded by a constant. One point of cutwidth, like pathwidth and treewidth, is that bounding it can ensure the tractability of some problems on graphs: when the input graph is assumed to have constant cutwidth, then we can design dynamic programming algorithms on the input graph.
In this paper, we study how to show upper bounds on the treewidth of graphs via the approach of partitioning its vertices into classes, and studying on the one hand the cutwidth of each class, and on the other hand the cutwidth of the quotient obtained by merging each class to a single vertex. As the quotient may in general feature parallel edges, and cutwidth accounts for these parallel edges, then we define the quotient as a multigraph, and work with multigraphs throughout the paper. Let us assume that we can show an upper bound of $x$ on the cutwidth of the quotient multigraph, and that we can show an upper bound of $y$ on the maximal cutwidth of a class of partition. We would then want to achieve a bound on the cutwidth of the original graph, as a function of $x$ and $y$. Intuitively, we want to do so by ordering the classes according to the order that achieves cutwidth $x$ on the quotient multigraph, and then ordering each class according to the order that achieves cutwidth at most $y$ within each class.
This technique of bounding cutwidth via vertex partitions has been studied earlier. barth1995bandwidth have studied how to bound the bandwidth or cutwidth of a graph as a function of the same measure on the quotient graph and on the classes of the partition; but their bounds depend on the degree of the graph. More recently, FeldmannM23 have applied the method in the specific case where the input graph is directed, and the classes of the partition are the strongly connected components (SCCs): the quotient multigraph is then called the condensation multigraph. Their work (FeldmannM23, Lemma 2.2) claims an upper bound of $x+y$ on the cutwidth of the graph $G$, again as a function of the (undirected) cutwidth $x$ of the condensation multigraph of $G$ and of the maximal (undirected) cutwidth $y$ of an SCC of $G$. This standalone lemma of FeldmannM23 is of independent interest, and we recently applied it to our study of the edge-minimum walk of modular length problem in amarilli2025edge.
In this paper, we specifically study this question of bounding the cutwidth of a graph using vertex partitions, namely, as a function of the cutwidth $x$ of the quotient multigraph and of the maximal cutwidth $y$ of a class. We show that the proof of (FeldmannM23, Lemma 2.2) contains an oversight: as we explain, their technique shows an upper bound of $2x+y$ and not $x+y$. Intuitively, the issue is that the order defined on the initial graph $G$, during its enumeration of the vertices of a class $C$, may cut some edges connecting $C$ to both preceding classes and succeeding classes in the ordering on the quotient multigraph: this means that we may cut more edges than the maximal number $x$ that we achieve on the quotient multigraph when we enumerate all vertices of $C$ at once.
This bounds of $2x+y$ leaves open the question of whether we can improve it, in particular to reach the bound of $x+y$ claimed in FeldmannM23. Our main technical contribution is to show (in Theorem 4.1) a bound of $1.5x+y$. This bound is obtained via a more careful choice of ordering on the initial graph $G$: we still order the classes of the partition in the order that achieves the cutwidth of $x$ on the quotient multigraph, but for each class of $G$ we choose among a minimum-cutwidth ordering (achieving cutwidth at most $y$) and its reverse (achieving the same cutwidth), intuitively depending on which one least worsens the cutwidth bound on the quotient multigraph.
Our result achieves a cutwidth bound $1.5x+y$ and not $x+y$, but as we show, this is unavoidable (and the bound of $x+y$ claimed in (FeldmannM23, Lemma 2.2) is in fact not correct). Specifically, we show in Proposition 5.1 that, for arbitrarily large $x$ and $y$, there are graphs whose cutwidth is $1.5x+y$: this implies that the bound of $1.5x+y$ cannot be improved in general.
Paper structure.
We give preliminaries in Section 2. We show our upper bound of $1.5x+y$ (Theorem 4.1) in Section 4, and show our lower bound (Proposition 5.1) in Section 5. We conclude in Section 6.
Acknowledgements.
We are grateful to the authors of FeldmannM23 for confirming the issue with (FeldmannM23, Lemma 2.2).
2 Preliminaries
[…]
We work with undirected and directed graphs and also with undirected and directed multigraphs, in which edges have multiplicities. We sometimes talk of a simple graph to mean a graph which is not a multigraph, i.e., the special case of a multigraph where no edge has multiplicity greater than $1$. Formally, we represent multigraphs using multisets of edges, where a multiset is a set in which some elements can be repeated (or, equivalently, have a multiplicity, which is a positive integer). We write $|S|$ to denote the cardinality of a multiset $S$ (summing over all multiplicities). We use double curly braces for multisets and for multiset comprehension notation, e.g., for $S$ a multiset, if we write $S^{\prime}\coloneq\{\{(x,y)\mid x,y\in S\}\}$, then we mean that $S^{\prime}$ is the multiset of ordered pairs of elements of $S$, taking each element of $S$ with its multiplicity (so that $|S^{\prime}|=(|S|)^{2}$). For $S_{1}$ and $S_{2}$ two multisets, we write $S_{1}\subseteq S_{2}$ to mean that the elements of $S_{1}$ are a subset of that of $S_{2}$ and the multiplicity of each element of $S_{1}$ is no greater than its multiplicity in $S_{2}$.
An undirected multigraph $(V,E)$ consists of a set $V$ of vertices and a multiset $E$ of undirected edges (i.e., pairs of vertices of $V$), and a directed multigraph $(V,E)$ consists of a set $V$ of vertices and a multiset $E$ of directed edges of the form $(u,v)$ with $u\neq v$ for $u,v\in V$. Note that we never allow self-loops in graphs. Given a directed multigraph $G=(V,E)$, the underlying undirected multigraph $G^{\prime}=(V,E^{\prime})$ of $G$ is obtained by forgetting the orientation of edges, i.e., we set $E^{\prime}\coloneq\{\{\{u,v\}\mid(u,v)\in E\}\}$. Note that $G^{\prime}$ may be a multigraph even when $G$ is a simple graph, e.g., if $E=\{(1,2),(2,1)\}$ then $E^{\prime}=\{\{\{1,2\},\{1,2\}\}\}$.
Cutwidth.
We consider the measure of cutwidth, defined on undirected graphs and multigraphs. We will also consider the cutwidth of directed graphs and multigraphs, but we always define the cutwidth of such graphs to be that of their underlying undirected multigraphs (i.e., we never consider definitions of cutwidth for directed graphs, such as the ones of chudnovsky2012tournament). Given a undirected (multi)graph $G=(V,E)$, a cut $V_{-},V_{+}$ of $V$ is a partition of $V$ into two sets $V_{-}$ and $V_{+}$, i.e., $V=V_{-}\sqcup V_{+}$ where $\sqcup$ denotes disjoint union. We say that an edge $e\in E$ crosses the cut $V_{-},V_{+}$ if it intersects both sides, namely, $e\cap V_{-}\neq\emptyset$ and $e\cap V_{+}\neq\emptyset$. Then the cutwidth of the cut $V_{-},V_{+}$ is the number of edges that cross the cut, counted together with their multiplicity, i.e., it is |$|\{\{e\in E\mid e\cap V_{-}\neq\emptyset\text{ and }e\cap V_{+}\neq\emptyset\}\}|$. We then define an ordering of $G$ as a strict total order $<$ on $V$, and say that a cut $V_{-},V_{+}$ of $V$ respects $<$ if we have $v_{-}<v_{+}$ for each $(v_{-},v_{+})\in V_{-}\times V_{+}$. The cutwidth of the ordering $<$ is then the maximum cutwidth of a cut that respects $<$. Note that, by symmetry, the cutwidth of an ordering $<$ is always the same as that of the reverse ordering $>$ defined by $x>y$ iff $y<x$. Finally, the cutwidth of $G$ is defined as the minimum cutwidth of an ordering of $G$.
Subdivisions.
Letting $G=(V,E)$ be an undirected multigraph, and $e=\{u,v\}$ be (one occurrence of) an edge, a subdivision of $G$ on $e$ is a multigraph obtained by removing (one occurrence of) $e$ from $E$ and replacing it by a path of length 2 that goes via a fresh intermediary vertex. Formally, it is an undirected multigraph $(V^{\prime},E^{\prime})$ obtained by letting $V^{\prime}\coloneq V\sqcup\{w\}$ where $w$ is a fresh vertex, and letting $E^{\prime}$ be constructed from $E$ by removing (one occurrence of) $e$ and adding the edges $\{u,w\}$ and $\{w,u\}$. Note that, if we take a multigraph and perform subdivision on every edge occurrence, then the result is always a simple graph. More generally, we define a multiedge-subdivision of $G$ as a graph $G^{\prime}=(V^{\prime},E^{\prime})$ as above, but where we consider any number $m$ of occurrences of $e$, with $m$ is at least $1$ and at most the number of occurrences of $e$ in $E$, and replace them with $m$ occurrences of the edges $\{u,w\}$ and $\{w,u\}$ (all sharing the same intermediate vertex $w$). Note that for $m=1$ we recover the definition of subdivisions given earlier.
We will need the following result about cutwidth and (multiedge-) subdivisions. In the case of subdivisions, the results appears to be folklore and appears for instance in MakedonS89. We provide a self-contained proof in Appendix 3 to show that the result also holds for multiedge-subdivisions.
{toappendix}
3 Proof of Section 3
{propositionrep}
Let $G=(V,E)$ be a multigraph, and let $G^{\prime}$ be obtained from $G$ by a multiedge-subdivision. Then the cutwidth of $G^{\prime}$ is equal to the cutwidth of $G$.
Proof.
Let $k$ be the cutwidth of $G$, and let $k^{\prime}$ be that of $G^{\prime}$. Let $w$ be the fresh vertex introduced by the multiedge-subdivision, $e=\{u,v\}$ the (multi)edge to which we apply the subdivision, $M$ the multiplicity of $e$ in $G$, and $1\leq m\leq M$ the number of occurrences of $e$ that we subdivide: we replace them in $G^{\prime}$ by $M$ occurrences of $e_{u}\coloneq\{u,w\}$ and $e_{v}\coloneq\{v,w\}$. To prove that $k=k^{\prime}$, we will first show that $k^{\prime}\leq k$ by inserting $w$ at an arbitrary position between $u$ and $v$ in an ordering $<$ of $G$ having cutwidth $k$: we will show that the resulting ordering $<^{\prime}$ of $G^{\prime}$ has cutwidth $\leq k$. Reciprocally, given an ordering of $G^{\prime}$, we will consider the ordering induced on $G$ by omitting $w$, and show that $k\leq k^{\prime}$.
We first show the first direction ($k^{\prime}\leq k$). Given an ordering $<$ of $V$ which achieves cutwidth $k$ for $G$, we assume up to reversing $<$ that $u$ occurs before $v$ in $<$. We then build an ordering ${<^{\prime}}$ of $V^{\prime}$ by inserting the fresh vertex $w$ at an arbitrary position between $u$ and $v$. For any cut $V_{-}^{\prime},V_{+}^{\prime}$ of $<^{\prime}$, considering the corresponding cut $V_{-},V_{+}$ of $<$, every edge crossing the cut $V_{-}^{\prime},V_{+}^{\prime}$ in $G^{\prime}$ except the two fresh edges $e_{u}$ and $e_{v}$ also crosses the cut $V_{-},V_{+}$. For the two fresh multiedges, we know that:
  • $e_{u}$ crosses the cut precisely when $u\in V_{-}^{\prime}$ but $w\in V_{+}^{\prime}$, and in this case the edge $e$ crosses the cut $V_{-},V_{+}$ because $v$ comes after $w$ in $<^{\prime}$ so $v\in V_{+}$ and of course $u\in V_{-}$;
  • $e_{v}$ crosses the cut precisely when $w\in V_{-}^{\prime}$ but $v\in V_{+}^{\prime}$, and in this case also the edge $e$ crosses the cut $V_{-},V_{+}$ because $u$ comes before $w$ in $<^{\prime}$ so $u\in V_{-}$ and $v\in V_{+}$.
It follows that only one of the multiedges $e_{u}$ or $e_{v}$ crosses the cut, and then the $m$ occurrences of that multiedge which are crossed in $<^{\prime}$ correspond to the $m$ additional occurrences of $e$ that are cut in $<$. Thus, the cutwidth of $V_{-}^{\prime},V_{+}^{\prime}$ in $G^{\prime}$ is no more than the cutwidth of $V_{-},V_{+}$ in $G$, so the cutwidth of ${<^{\prime}}$ is at most $k$, which establishes that $k^{\prime}\leq k$.
We then show the second direction ($k\leq k^{\prime}$). Given an ordering ${<^{\prime}}$ of $V^{\prime}$ which achieves cutwidth $k^{\prime}$ for $V^{\prime}$, we assume again up to reversing ${<^{\prime}}$ that $u$ occurs before $v$. We build the ordering $<$ of $V$ simply by removing the fresh vertex $w$. Consider any cut $V_{-},V_{+}$ of $<$, and consider some cut $V_{-}^{\prime},V_{+}^{\prime}$ of ${<^{\prime}}$ which is equal to $V_{-},V_{+}$ up to the fresh vertex, i.e., we have $V_{-}\subseteq V_{-}^{\prime}\subseteq V_{-}\sqcup\{w\}$ and likewise $V_{+}\subseteq V_{+}^{\prime}\subseteq V_{+}\sqcup\{w\}$. For any edge of $G$ which is cut by $V_{-},V_{+}$ and is different from $e$, we know that the same edge in $G^{\prime}$ is also cut in $V_{-}^{\prime},V_{+}^{\prime}$. Now, the edge $e$ crosses the cut precisely when $u\in V_{-}$ but $v\in V_{+}$. In this case, we have $u\in V_{-}^{\prime}$ and $v\in V_{+}^{\prime}$, so it must be the case that one of the edges $e_{u}$ and $e_{v}$ crosses the cut $V_{-}^{\prime},V_{+}^{\prime}$, which compensates the $m$ additional occurrences of $e$ in $G$ that are cut by $V_{-},V_{+}$. Hence, the cutwidth of $V_{-},V_{+}$ is no more than that of $V_{-}^{\prime},V_{+}^{\prime}$, so that $k\leq k^{\prime}$. This shows that $k=k^{\prime}$ and concludes the proof. ∎
Vertex partitionings.
We will bound the cutwidth of undirected (multi)graphs $G$ using vertex partitionings, which we now define. A vertex partitioning of a set $V$ of vertices is simply a partition $P=\{C_{1},\ldots,C_{k}\}$ of $V$: each class $C_{i}$ for $1\leq i\leq k$ is a nonempty subset of $V$, the classes are pairwise disjoint, and their union is $V$. Given a vertex partitioning of the vertices of an undirected (multi)graph $G=(V,E)$, we can distinguish two kinds of edges of $G$: the internal edges, where both endpoints belong to the same class; and the external edges, which connect endpoints belonging to different classes.
A subgraph of a (multi)graph $G=(V,E)$ is simply a (multi)graph $(V,E^{\prime})$ with $E^{\prime}\subseteq E$, and the subgraph induced by a subset $C\subseteq V$ of vertices is the subgraph $(V,E_{|C})$ where we keep the internal edges of $E$ for which both endpoints are in $C$ (with their multiplicities), formally, $E_{|C}\coloneq\{\{e\in E\mid e\subseteq C\}\}$. We will often abuse notation and identify the classes of $P$ with the subgraphs of $G$ that they induce.
From the undirected (multi)graph $G$ and the vertex partition $P=\{C_{1},\ldots,C_{k}\}$, in addition to the subgraphs induced by the $C_{i}$ for $1\leq i\leq k$, we will also consider the quotient multigraph where we intuitively merge each class of the partition into a single vertex, and only keep the external edges between these vertices. Formally, let $\phi_{P}\colon V\to P$ be the function that maps each vertex of $V$ to its class in $P$. The quotient multigraph $G/P$ of $G$ under $P$ is then the undirected multigraph with vertex set $P$, and with the multiset of edges $\{\{\{\phi_{P}(u),\phi_{P}(v)\}\mid\{u,v\}\in E\text{\leavevmode\nobreak\ s.t.\leavevmode\nobreak\ }\phi_{P}(u)\neq\phi_{P}(v)\}\}$. Note that the quotient multigraph is in general a multigraph even if $G$ is a simple graph, e.g., if $G$ contains the edges $\{1,3\}$ and $\{2,3\}$ and $P$ contains the classes $\{1,2\}$ and $\{3\}$.
We also define quotient graphs over directed (multi)graphs in the expected way, in particular for the case of SCCs that we define next. Formally, for $G=(V,E)$ a directed (multi)graph, for $P$ a vertex partitioning of $V$, and for $\phi_{P}\colon V\to P$ the function mapping the vertices of $V$ to their classes in $P$, the quotient multigraph of $G$ by $P$ is the directed multigraph $G/P$ whose vertex set is $P$ and whose multiset of edges is $\{\{(\phi_{P}(u),\phi_{P}(v))\mid(u,v)\in E\}\}$.
SCCs.
One special case of vertex partitioning which we will study is the well-known decomposition of directed (multi)graphs into Strongly Connected Components (SCCs). Specifically, we say that two vertices $u$ and $v$ of a directed (multi)graph $G=(V,E)$ are in the same SCC if there is a directed path from $u$ to $v$ in $G$ and a directed path from $v$ to $u$ in $G$: this relation is clearly an equivalence relation, so it defines a vertex partitioning $P$. Note that, for each class $C$ of $P$, the subgraphs of $G$ induced by the classes of $P$ are exactly the strongly connected components of $G$ in the usual sense. Further, the quotient multigraph $G/P$ is then called the condensation multigraph $G^{\prime}$ of $G$, i.e., $G^{\prime}$ is the directed acyclic multigraph obtained by condensing the SCCs of $G$. Note that, in most of the works that consider the decomposition of a directed graph into SCCs, the condensation graph is defined as a directed acyclic simple graph over SCCs: but in this note we see $G^{\prime}$ as a (directed acyclic) multigraph (like in FeldmannM23), with the multiplicity of every edge from $C$ to $C^{\prime}$ in $G^{\prime}$ intuitively denoting how many edges of $G$ go from a vertex in $C$ to a vertex in $C^{\prime}$. This distinction will be important when studying the cutwidth of the condensation multigraph.
4 Upper bound
In this section, we show the following general result on vertex partitionings of undirected multigraphs:
Theorem 4.1.
Let $G=(V,E)$ be an undirected multigraph, $P$ a partition of $G$, and $G/P$ be the quotient multigraph of $G$ under $P$. If the cutwidth of $G/P$ is $x$ and the cutwidth of every class of $P$ is at most $y$, then the cutwidth of $G$ is at most $1.5x+y$.
This result holds for arbitrary vertex partitionings. If we instantiate it to the specific case of vertex partitioning defined by SCCs on directed graphs, we obtain the following corollary:
Corollary 4.2.
Let $G$ be a directed multigraph and $G^{\prime}$ be its condensation multigraph. If the cutwidth of $G^{\prime}$ is $x$ and the cutwidth of every SCC of $G$ is at most $y$, then the cutwidth of $G$ is at most $1.5x+y$.
Remember that, here, the cutwidths are always taken on the underlying undirected multigraphs. The statement of the corollary is similar to (FeldmannM23, Lemma 2.2), except that the bounds are worse ($1.5x+y$ instead of $x+y$). In the rest of this section, we prove Theorem 4.1. We first show a weaker bound of $2x+y$ with a simple approach that uses the ordering defined in FeldmannM23. We next show how a careful choice of ordering lowers the bound to $1.5x+y$. We will show in the next section that this bound cannot be improved in general.
Simple bound in $\bm{2x+y}$.
Let $G=(V,E)$ be an undirected multigraph and $P$ be a vertex partitioning of $G$. Remember that we identify each class $C$ of $P$ with the subgraph of $G$ that it induces. We will define an ordering of $G$, and we will require that it is defined by first choosing an ordering on the classes of $P$, and then ordering the vertices inside each class $C$ according to some ordering on $C$. Following the terminology of (barth1995bandwidth, Section 2.3), we say that such an ordering of $G$ is compatible with $P$: we will only consider such orderings in the proof.
For each class $C\in P$, we let $<_{C}$ be an ordering on $C$ that achieves a cutwidth of at most $y$ on $C$. Further, let $<_{G/P}$ be an ordering on the quotient multigraph $G/P$ that achieves the optimal cutwidth of $x$ on $G/P$. We now define a compatible ordering $<$ of $G$ like in FeldmannM23: we first order vertices following the order of classes given by $<_{G/P}$, and then inside each class $C$ we order the vertices according to the ordering $<_{C}$. Formally, for $u,v\in V$, letting $C_{i}$ and $C_{j}$ be the respective classes of $u$ and $v$, we set $u<v$ if we have $C_{i}<_{G/P}C_{j}$ or if we have $C_{i}=C_{j}$ and $u<_{C}v$ for $C_{i}=C=C_{j}$.
Let $C$ be a class, and let $V_{-}^{\prime}$ (resp. $V_{+}^{\prime}$) consist of all vertices of classes before $C$ (resp. after $C$) in $<_{G/P}$. Any cut $C_{-},C_{+}$ of $C$ induces a cut $V_{-}^{\prime}\sqcup C_{-},C_{+}\sqcup V_{+}^{\prime}$ on $G$ that respects the ordering $<$. Let $\Xi$ be the multiset of edges of $E$ that cross the cut. We can partition $\Xi$ into the following disjoint sets:
$\Xi=\Xi_{C,C}\sqcup\Xi_{V_{-}^{\prime},C}\sqcup\Xi_{C,V_{+}^{\prime}}\sqcup\Xi_{V_{-}^{\prime},V_{+}^{\prime}},$
where $\Xi_{X,Y}$ is the set of edges crossing the cut with one extremity in $X$ and the other in $Y$. For instance, $\Xi_{C,C}$ is the subset of the edges internal to $C$ which cross the cut, whereas $\Xi_{V_{-}^{\prime},C}$ are the edges with one extremity in $V_{-}^{\prime}$ and the other extremity in $C$ which cross the cut (and therefore $\Xi_{V_{-}^{\prime},C}=\Xi_{V_{-}^{\prime},C_{+}}$).
The number of edges $|\Xi_{C,C}|$ is bounded by $y$ because the restriction of $<$ to $C$ achieves a cutwidth of at most $y$, and the quantities $|\Xi_{V_{-}^{\prime},C}|+|\Xi_{V_{-}^{\prime},V_{+}^{\prime}}|$ and $|\Xi_{C,V_{+}^{\prime}}|+|\Xi_{V_{-}^{\prime},V_{+}^{\prime}}|$ are each bounded by $x$ by considering the order $<_{G/P}$ on the quotient multigraph. Hence, the number of edges crossing the cut is at most $|\Xi_{V_{-}^{\prime},C}|+|\Xi_{C,V_{+}^{\prime}}|+|\Xi_{V_{-}^{\prime},V_{+}^{\prime}}|+|\Xi_{C,C}|\leq 2x+y$. We have thus proved that ordering $<$ has cutwidth at most $2x+y$.
More elaborate bound in $\bm{1.5x+y}$.
To show the claimed bound of $1.5x+y$, we now define the ordering $<$ more carefully. We will still order vertices first according to the order $<_{G/P}$, so that in particular the order is still compatible: but then within each class $C$ we will choose between the ordering $<_{C}$ and its reverse $>_{C}$.
Let $C$ be a class, and let $V_{-}^{\prime}$ (resp. $V_{+}^{\prime}$) consist of all vertices of classes before $C$ (resp. after $C$) in $<_{G/P}$. As we explained previously, every cut $C_{-},C_{+}$ of $C$ according to $<_{C}$ induces a cut $V_{-}^{\prime}\sqcup C_{-},C_{+}\sqcup V_{+}^{\prime}$ on $G$. The same split of $C$ into $C_{+},C_{-}$ is also a cut of $C$ according to $>_{C}$ where the vertices of $C_{+}$ now come before those of $C_{-}$ and this cut according to $>_{C}$ induces a cut $V_{-}^{\prime}\sqcup C_{+},C_{-}\sqcup V_{+}^{\prime}$ on $G$.
We can partition the edges of $G$ as follows, where $E_{X,Y}$ denotes the edges having an extremity in $X$ and the other in $Y$. The partitioning is similar to the one above, except that we partition all the edges and not only those crossing the cut of $C$ according to $<_{C}$, and we distinguish more edge types:
$E=E_{V_{-}^{\prime},V_{-}^{\prime}}\sqcup E_{V_{+}^{\prime},V_{+}^{\prime}}\sqcup E_{C,C}\sqcup E_{V_{-}^{\prime},C_{-}}\sqcup E_{C_{+},V_{+}^{\prime}}\sqcup E_{V_{-}^{\prime},C_{+}}\sqcup E_{C_{-},V_{+}^{\prime}}\sqcup E_{V_{-}^{\prime},V_{+}^{\prime}}$ (1)
  • The edges from $E_{V_{-}^{\prime},V_{-}^{\prime}}\sqcup E_{V_{+}^{\prime},V_{+}^{\prime}}$ (internal to other classes than $C$, or connecting classes located on the same side of the cut): these edges will never cross the cut (no matter whether we order vertices of the various classes $C^{\prime}$ according to $<_{C^{\prime}}$ or according to $>_{C^{\prime}}$).
  • The edges from $E_{C,C}$ are internal to $C$: at most $y$ of these cross the cut, because we will order the vertices of $C$ either according to $<_{C}$ or according to $>_{C}$ (and both achieve the same cutwidth).
  • The edges from $E_{V_{-}^{\prime},C_{+}}\sqcup E_{C_{-},V_{+}^{\prime}}$ cross the cut $V_{-}^{\prime}\sqcup C_{-},C_{+}\sqcup V_{+}^{\prime}$ according to $<_{C}$ (but not the cut $V_{-}^{\prime}\sqcup C_{+},C_{-}\sqcup V_{+}^{\prime}$ according to $>_{C}$).
  • The edges from $E_{V_{-}^{\prime},C_{-}}\sqcup E_{C_{+},V_{+}^{\prime}}$ cross the cut $V_{-}^{\prime}\sqcup C_{+},C_{-}\sqcup V_{+}^{\prime}$ according to $>_{C}$ (but not the cut $V_{-}^{\prime}\sqcup C_{-},C_{+}\sqcup V_{+}^{\prime}$ according to $<_{C}$).
  • The edges from $E_{V_{-}^{\prime},V_{+}^{\prime}}$ always cross the cut (no matter whether we order the vertices of the classes $C^{\prime}$ according to $<_{C^{\prime}}$ or according to $>_{C^{\prime}}$).
Figure 1 illustrates all the edge types which may cross the cut. We make the following claim, which implies that one of $<_{C}$ and its reverse will achieve the cutwidth bound:
Claim 1.
Let $<_{G/P}$ be an ordering on $G/P$ that achieves the optimal cutwidth $x$ on $G/P$. Let $V_{-}^{\prime}$ (resp. $V_{+}^{\prime}$) consist of all vertices of classes before (resp. after) $C$ in $<_{G/P}$. For every class $C$ and ordering $<_{C}$ on $C$, at least one of the following claims is true:
  • For every cut $C_{-},C_{+}$ of $C$ according to $<_{C}$, we have $|E_{V_{-}^{\prime},C_{+}}|+|E_{C_{-},V_{+}^{\prime}}|+|E_{V_{-}^{\prime},V_{+}^{\prime}}|\leq 1.5x$
  • For every cut $C_{-},C_{+}$ of $C$ according to $<_{C}$, we have $|E_{V_{-}^{\prime},C_{-}}|+|E_{C_{+},V_{+}^{\prime}}|+|E_{V_{-}^{\prime},V_{+}^{\prime}}|\leq 1.5x$
This claim suffices to conclude the proof of Theorem 4.1. Indeed, the first (resp. second) case in the claim bounds the number of external edges crossed in cuts compatible with $<_{C}$ (resp., with $>_{C}$). We can therefore use the claim on each class $C$ to choose to order its vertices either according to $<_{C}$ or according to $>_{C}$. Overall, this gives a compatible order $<$, of which we can show that it has cutwidth at most $1.5x+y$. Indeed, let us consider a cut $V_{-},V_{+}$. If $V_{+}$ is empty no edge can cross the cut, so we may assume that $V^{+}$ is non-empty. Let $C$ be the first class in $<_{G/P}$ which contains a vertex from $V^{+}$. Then $V_{-},V_{+}$ is the cut induced by $C_{-},C_{+}$ where $C_{-}=V_{-}\cap C$ and $C_{+}=V_{+}\cap C$. We can then bound the number of external edges crossed in the cut $V_{-},V_{+}$ by $1.5x$ from the claim statement (because we chose to order $C$ according to $<_{C}$ or to $>_{C}$ depending on which of the two statements of the claim is true), and the number of internal edges crossed in the cut is bounded by the number of internal edges in $C$ crossed in the cut, which is at most $y$ because the order on $C$ is either $<_{C}$ or $>_{C}$. This allows us to conclude the proof of Theorem 4.1.
Hence, all that remains is to show Claim 1. The intuition is that we consider the worst cut according to the first equation: either the cut in question satisfies the bound of $1.5x$ and we use the first case, or it does not. In this second case, partitioning the external edges involving vertices of $C$ in 4 types according to that cut (like the four last sets of Equation 1), we show how the violation of the first case implies that the second case is satisfied by any cut of $C$. The formal proof follows:
Proof 4.3 (Proof of Claim 1).
Let us first observe that, given that the ordering $<_{G/P}$ achieves a cutwidth bound of at most $x$ on the quotient multigraph $G/P$, then by considering the cut right before $C$ (i.e., the cut $V_{-}^{\prime},C\sqcup V_{+}^{\prime}$) we know that $|E_{V_{-}^{\prime},C}|+|E_{V_{-}^{\prime},V_{+}^{\prime}}|\leq x$, where $E_{V_{-}^{\prime},C}$ denotes the external edges involving a vertex of $C$ and a vertex of a class before $C$ in $<_{G/P}$. Consequently, for any cut $C_{-},C_{+}$ of $C$ according to $<_{C}$, given that $E_{V_{-}^{\prime},C}=E_{V_{-}^{\prime},C_{-}}\sqcup E_{V_{-}^{\prime},C_{+}}$, we have:
$|E_{V_{-}^{\prime},C_{-}}|+|E_{V_{-}^{\prime},C_{+}}|+|E_{V_{-}^{\prime},V_{+}^{\prime}}|\leq x$ (2)
and similarly, considering the cut right after $C$, we obtain:
$|E_{C_{-},V_{+}^{\prime}}|+|E_{C_{+},V_{+}^{\prime}}|+|E_{V_{-}^{\prime},V_{+}^{\prime}}|\leq x$ (3)
These 2 cuts are depicted by dotted vertical lines on the left of Figure 1.
We now fix a cut $C_{-},C_{+}$ of $C$ according to $<_{C}$ with $C_{+}$ nonempty that maximizes the number of external edges crossing the cut when ordering $C$ according to $<_{C}$, i.e., which maximizes the value $n\coloneq|E_{V_{-}^{\prime},C_{+}}|+|E_{C_{-},V_{+}^{\prime}}|+|E_{V_{-}^{\prime},V_{+}^{\prime}}|$ in the left-hand-side of the inequality in the first case of Claim 1. There are two cases: either we have $n\leq 1.5x$, or we have $n>1.5x$. In the first case, we can conclude immediately that the first case of Claim 1 holds, because our choice for the cut $C_{-},C+$ makes $n$ an upper bound on the left-hand side of the first inequality. Hence, in the rest of the proof, we focus on the second case, where we have:
$|E_{V_{-}^{\prime},C_{+}}|+|E_{C_{-},V_{+}^{\prime}}|+|E_{V_{-}^{\prime},V_{+}^{\prime}}|>1.5x$ (4)
[…]
$|E_{V_{-}^{\prime},C_{-}}|+|E_{C_{+},V_{+}^{\prime}}|<0.5x$ (5)
which implies in particular the following:
$|E_{V_{-}^{\prime},C_{-}}|<0.5x$ (6)
$|E_{C_{+},V_{+}^{\prime}}|<0.5x$ (7)
We will use these properties of the fixed cut $C_{-},C_{+}$ to show that the second case of Claim 1 holds. Let us therefore show that, for every cut $C_{-}^{\prime},C_{+}^{\prime}$ of $C$ according to $<_{C}$, we can bound the cutwidth of the cut $C_{+}^{\prime},C_{-}^{\prime}$ for $>_{C}$, i.e., we have the bound $|E_{V_{-}^{\prime},C_{-}^{\prime}}|+|E_{C_{+}^{\prime},V_{+}^{\prime}}|+|E_{V_{-}^{\prime},V_{+}^{\prime}}|\leq 1.5x$, which is the inequality in the second case of Claim 1. To do so, we will use partition the external edges involving $C$ in 4 types according to the fixed cut $C_{-},C_{+}$ chosen earlier.
We distinguish two cases:
  • The cut $C_{-}^{\prime},C_{+}^{\prime}$ is “before” $C_{-},C_{+}$, i.e., we have $C_{-}^{\prime}\subseteq C_{-}$ and $C_{+}^{\prime}\supseteq C_{+}$. In this case, we have $E_{V_{-}^{\prime},C_{-}^{\prime}}\subseteq E_{V_{-}^{\prime},C_{-}}$, and we always have $E_{C_{+}^{\prime},V_{+}^{\prime}}\subseteq E_{C_{+},V_{+}^{\prime}}\sqcup E_{C_{-},V_{+}^{\prime}}$. Thus, we have:
    $|E_{V_{-}^{\prime},C_{-}^{\prime}}|+|E_{C_{+}^{\prime},V_{+}^{\prime}}|+|E_{V_{-}^{\prime},V_{+}^{\prime}}|\leq|E_{V_{-}^{\prime},C_{-}}|+|E_{C_{+},V_{+}^{\prime}}|+|E_{C_{-},V_{+}^{\prime}}|+|E_{V_{-}^{\prime},V_{+}^{\prime}}|$
[…]
  • The cut $C_{-}^{\prime},C_{+}^{\prime}$ is “after” $C_{-},C_{+}$, i.e., we have $C_{+}^{\prime}\subseteq C_{+}$ and $C_{-}^{\prime}\supseteq C_{-}$. In this case, we have $E_{C_{+}^{\prime},V_{+}^{\prime}}\subseteq E_{C_{+},V_{+}^{\prime}}$, and we always have $E_{V_{-}^{\prime},C_{-}^{\prime}}\subseteq E_{V_{-}^{\prime},C_{-}}\sqcup E_{V_{-}^{\prime},C_{+}}$. Thus:
    $|E_{V_{-}^{\prime},C_{-}^{\prime}}|+|E_{C_{+}^{\prime},V_{+}^{\prime}}|+|E_{V_{-}^{\prime},V_{+}^{\prime}}|\leq|E_{V_{-}^{\prime},C_{-}}|+|E_{V_{-}^{\prime},C_{+}}|+|E_{C_{+},V_{+}^{\prime}}|+|E_{V_{-}^{\prime},V_{+}^{\prime}}|$
[…]
Hence, the second case of Claim 1 then holds. This concludes the proof.
[…]
Figure 1: Partitioning edges that may contribute to the cuts before or after reversing $<_{C}$.
5 Lower bound
In this section, we show that the constant factor of $1.5$ in the statement of Theorem 4.1 cannot be improved: we show in particular that the upper bound of $x+y$ claimed by FeldmannM23 does not hold. Here is the formal statement that we show:
Proposition 5.1.
For each even integer $x\geq 2$, for each integer $y\geq 1.5x$, there is a directed graph $H$ whose cutwidth is $1.5x+y$, where $x$ is the cutwidth of the condensation multigraph of $H$, and $y$ is the maximum cutwidth of an SCC from $H$.
We show in Figure 2 a directed graph obtained in our proof of Proposition 5.1 for the case where $x=2$ and $y=3$. This directed graph has cutwidth 6 even though $x+y=5$.
[…]
Claim 2.
The cutwidth of $G$ is $\min(1.5x+y,\max(2y,x))$.
Let us first explain why the claim suffices to prove Proposition 5.1. Let $x\geq 2$ be even and let $y\geq 1.5x$, and let us build the multigraph $G$ for these values of $x$ and $y$: by Claim 2, we know that $G$ has cutwidth $1.5x+y$, because $y\geq 1.5x$ ensures that $2y\geq x$ and also ensures that $2y\geq 1.5x+y$, so the expression of the claim evaluates to $1.5x+y$. Now, let $G^{\prime}$ be the undirected graph obtained from $G$ by subdividing each occurrence of each edge: by repeated application of Proposition 3, we know that $G^{\prime}$ has the same cutwidth as $G$, and $G^{\prime}$ is now a simple undirected graph. Last, let $H$ be the directed graph obtained by orienting the edges of $G^{\prime}$ so that the only non-trivial connected component is the one featuring the vertices $\{2,3,4\}$ along with the intermediate vertices added in the subdivision: as $y\geq 1$, we can achieve this by picking different orientations for two of the paths of length $2$ which connect $2$ and $3$ in the subdivision, doing the same for the paths of length $2$ connecting $3$ and $4$, and ordering all other edges along a topological sort to ensure that no other non-trivial SCCs exist. Figure 2 shows the graph $H$ obtained at the end of the process (in the case where $x=2$ and $y=3$).
The end result is a directed graph $H$ whose condensation multigraph has the same cutwidth as the quotient multigraph of $G$ by $P$ (hence, $x$) and where the cutwidth of SCCs is at most $y$, but the cutwidth of $H$ is the same as that of $G$, i.e., it is $1.5x+y$. This suffices to show Proposition 5.1.
11115555222233334444111155552222
Figure 2: Example of the directed graph $H$ obtained in the proof of Proposition 5.1 (left), together with its condensation multigraph (right). For this graph, the maximum cutwidth of an SCC is $y=3$ (indeed the only nontrivial SCC is formed by the filled brown nodes ($\{2,3,4\}$ and the nodes connecting them), connected by dashed brown edges), and the cutwidth of the condensation multigraph is $x=2$, however it can be checked that the cutwidth of $H$ is $61.5x+y=$ 6 and not $x+y=5$.
Hence, all that remains is to show the claim:
Proof 5.2 (Proof of Claim 2).
First, we simplify the graph. To simplify the proof, we observe that by Section 3, $G$ has the same cutwidth as the multigraph $K$ obtained by removing node $5$ and replacing its incident edges by $x/2$ edges connecting $2$ to $4$. Consequently, we want to show that $K$ has the cutwidth stated by the claim.
In $K$, ordering $2,3,1,4$ achieves a cutwidth of $1.5x+y$, whereas the number of edges that cross the successive cuts of ordering $2,4,3,1$ are respectively $y+x/2$, $2y$, and $x$ so that this ordering has cutwidth $\max(2y,x)$. (It cannot be the case that the maximum is achieved by $y+x/2$ because this would imply $y+x/2>2y$ so that $x/2>y$ but then $y+x/2<x$.) This proves that the cutwidth of $G$ is at most $\min(1.5x+y,\max(2y,x))$. (Note that the upper bound of $1.5x+y$ could alternatively be viewed as a consequence of Theorem 4.1.)
Let us now prove a corresponding lower bound on the cutwidth of $K$. Let $C$ denote the subset $\{2,3,4\}$ of vertices of $K$, and let $<_{K}$ denote an ordering on the $4$ nodes of $K$. Consider first the case where the ordering is monotone on the identifiers of nodes from $C$. Up to reversing the ordering (which does not change the cutwidth), we may assume $42<_{K}3<_{K}$ 4. If $31<_{K}$ 3 then $K$ contains $1.5x+y$ edges crossing the cut ($(\{2,1\},\{3,4\})$. If $31>_{K}$ 3, then $K$ contains $1.5x+y$ edges crossing the cut ($(\{2,3\},\{1,4\})$. This shows that such monotone orders have cutwidth at least $1.5x+y$.
Consider now the case where the ordering $<_{K}$ is non-monotone. Up to reversing the order, and because 2 and 4 play symmetric roles, we can assume that the ordering $<_{K}$ is $32<_{K}4<_{K}$ 3. Now, any cut of $K$ separating $\{2,4\}$ from $\{3\}$ is crossed by $2y$ edges, whereas any cut separating $1$ from $3$ is crossed by $x$ edges. This shows that such non-monotone orders have cutwidth at least $\max(2y,x)$.
As a consequence, $K$ and $G$ have cutwidth $\min(1.5x+y,\max(2y,x))$, which concludes the proof of the claim.
[…]
Figure 3: Multigraph $G$ used in the proof of Proposition 5.1 (left) and its condensation multigraph (right).
Let us last remark that the lower bound shown in Proposition 5.1 holds for the specific graphs that we construct, but not for general graphs. More precisely, consider a directed graph $H$ and its condensation multigraph $H^{\prime}$. The cutwidth of $H$ cannot be less than the maximal cutwidth $y$ of an SCC of $G$, because SCCs are subgraphs of $G$ and cutwidth is monotone under taking subgraphs. However, the cutwidth $x$ of $H^{\prime}$ is not a lower bound on the cutwidth of $G$. Indeed:
Proposition 5.3.
For each $n\geq 3$, we can build a directed graph $G_{n}$ where the SCC condensation multigraph has cutwidth $x=n$, the maximal cutwidth of an SCC is $2$, but the graph $G_{n}$ itself has constant cutwidth.
[…]
Figure 4: Graph $G_{n}$ used in the proof of Proposition 5.3 (left), and its condensation multigraph (right).
Proof 5.4.
We picture the graph $G_{n}$ used to prove the result in Figure 4 (left).
The graph $G_{n}=(V_{n},E_{n})$ has vertices $V_{n}=\{u_{1},\ldots,u_{n}\}\sqcup\{v_{1},\ldots,v_{n}\}$ and has as edges $E_{n}$ two directed cycles $(u_{1},u_{2}),\ldots,(u_{n-1},u_{n}),(u_{n},u_{1})$ and $(v_{1},v_{2}),\ldots,(v_{n-1},v_{n}),(v_{n},v_{1})$, along with $n$ edges $(u_{i},v_{i})$ for each $1\leq i\leq n$. Thus, the SCCs of $G_{n}$ are $\{u_{1},\ldots,u_{n}\}$ and $\{v_{1},\ldots,v_{n}\}$, each of which is a cycle with cutwidth $2$.
The condensation multigraph multigraph (pictured in Figure 4, right) consists of two vertices connected by $n$ parallel edges, so it has cutwidth $n$. However, the ordering $u_{1}<v_{1}<\cdots<u_{n}<v_{n}$ achieves a cutwidth of at most 5: in each cut we have at most 2 internal edges cut in each of the two SCCs, and at most one external edge cut. This concludes the proof.
6 Conclusion and Future Work
We have shown that, for any graph $G$ and vertex partitioning of $G$, if we let $x$ be the cutwidth of the quotient multigraph of $G$ and let $y$ be the maximal cutwidth of a class of $G$, then the cutwidth of $G$ can be bounded with $1.5x+y$; and there are arbitrarily large values of $x$ and $y$ for which this bound cannot be improved.
Our work also leaves open the question of whether better bounds can be shown, in particular whether the upper bound of Theorem 4.1 can be improved in some settings. In particular, our lower bound of Proposition 5.1 is only shown in the case where $y\geq 1.5x$, leaving open the question of whether better upper bounds than Theorem 4.1 can be shown in the setting where $y<1.5x$. Another more general question for future research is whether the vertex partitioning approach that we studied could also be applied to other width measures beyond cutwidth.